# Global-Local Progressive Integration Network for Blind Image Quality Assessment

### v1

@@ V-D2 The proposed dataset
VI Conclusion
In this study, we have enhanced the effectiveness and robustness of BIQA by optimizing the network architecture and augmenting the training data. We proposed a novel BIQA model GlintIQA that leverages the synergy between CNNs and ViTs to extract complementary local and global multi-grained features. This hybrid architecture mitigates patch-level information loss and inductive bias limitations inherent in ViTs. GlintIQA employs a progressive integration strategy with interactive stacking channel-wise self-attention and spatial enhancement modules, efficiently handling diverse features across both channels and spatial dimensions. Furthermore, to overcome the issue of limited content variation in synthetic distortion datasets, we propose a novel content-guided method for automatic image quality labeling based on subjective quality scores. The proposed content-guided labeling method plays a pivotal role in significantly diversifying and enriching the training data, resulting in a substantial boost in model generalization ability. The innovative architecture of GlintIQA, combined with the proposed content-driven dataset, leads to a significant improvement in BIQA performance, as evidenced by extensive single-dataset and cross-dataset evaluations.

### v2

@@ IV-D3 The proposed SAQT-IQA dataset
V Conclusion
This study have addressed key limitations in NR-IQA by focusing on advancements in both model architecture and training data. We proposed GlintIQA, a novel NR-IQA model that leverages the synergy between CNNs and ViTs to extract complementary local and global multi-grained features. To effectively leverage these features, we proposed a progressive integration strategy that employs cascaded channel-wise self-attention and spatial interaction enhancement modules to dynamically weight feature channels and enhance spatial interactions, leading to improved quality assessment accuracy. To address limited content diversity in synthetic distortion datasets, we developed the semantic-aligned quality transfer method that enables the creation of the SAQT-IQA dataset, This dataset effectively mitigates data scarcity limitations and enhances model generalization. The collective impact of the GlintIQA architecture and the SAQT-IQA dataset is a demonstrably superior NR-IQA performance, confirmed through extensive experimental validation across diverse datasets.